### v2

@@ 3.2 Assumptions and Main Identification Result
Assumption 1 (Model).
Each coordinate of $g\left(Z_{it},A_{i},\eta_{it}\right)$ depends on one coordinate of $\eta_{it}$ strictly monotonically.
[…]
Assumption 1 requires justification when $d_{X}>1$, $d_{\eta}>1$, and each endogenous coordinate of $X_{it}$ admits a different coordinate of $\eta_{it}$. For example, it is possible that firms have different departments making hiring and investment decisions separately. Such decisions are based on different signals from the labor market and financial market. Therefore, it requires different coordinates of $\eta_{it}$ to enter the corresponding coordinate of $X_{it}$. Additionally, it is natural to allow firms to take advantage of positive shocks from labor market and financial market by hiring more employees and making more investments, leading to coordinate-wise strict monotonicity of $g\left(\cdot\right)$ in $\eta_{it}$. As a result, Assumption 1 is satisfied in this context.
It is worth highlighting that Assumption 1 does not restrict how different coordinates of $\eta_{it}$ are correlated. Thus, without loss of generality, I can redefine $\eta_{it}$ such that $d_{\eta}=d_{X}$ and each coordinate of $g\left(\cdot\right)$ depends on the corresponding coordinate of $\eta_{it}$. Moreover, Assumption 1 does not restrict how $A_{i}$ and $\eta_{it}$ are correlated, which could be important to empirical applications. Still, Assumption 1 restricts how the vector of $\eta_{it}$ enters the $g\left(\cdot\right)$ function that determines the value of $X_{it}$, which is a limitation of the paper. In Section 3.3, I discuss two other specifications to allow multiple coordinates of $\eta_{it}$ to enter one coordinate of $g\left(\cdot\right)$ when additional timing assumptions are considered plausible by practitioners and show how to adapt my method to these alternative specifications.
@@
Let $W_{i}\coloneqq W\left({\bf X}_{i},{\bf Z}_{i}\right)$, where $W:\mathbb{R}^{T\times\left(d_{X}+d_{Z}\right)}\mapsto\mathbb{R}^{d_{W}}$ is known. Suppose $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$ for any fixed $t$.
Assumption 2 is an index exclusion restriction that requires the conditional distribution of $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)$ to depend only on $W_{i}$ which is a known function of ${\bf X}_{i}$ and ${\bf Z}_{i}$. It is similar to Assumption 2.1 of Altonji and Matzkin (2005) and Assumption A2 of Liu, Poirier, and Shiu (2024). A similar condition is also imposed by Wooldridge (2019). Liu, Poirier, and Shiu (2024) prove a non-identification result when the distribution of $\left.A_{i}\right|{\bf X}_{i}$ is left unrestricted in a nonlinear panel data model and suggest that restrictions on $A_{i}$ are generally required to identify the APE and LAR. Given $W_{i}$, the distribution of $A_{i}$ no longer depends on $X_{it}$ and $Z_{it}$ for any fixed $t$. In other words, $W_{i}$ is sufficient for $A_{i}$. Assumption 2 is used to (i) construct the conditional control variable for $\eta_{it}$, which I discuss in detail in relation to Assumption 3; and (ii) to control for $A_{i}$ when using the LIE argument to exclude $X_{it}$ from the conditioning set of $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ in (3.4), which I present in the proof of Theorem 1.
As pointed out by Liu, Poirier, and Shiu (2024), there are several ways to justify Assumption 2 from the literature. I follow their discussion to provide three justifications of Assumption 2. The first two methods are essentially parametric, resulting in a simpler expression for $W_{i}$, while the last one is nonparametric, leading to a more complex expression for $W_{i}$. This distinction highlights the trade-off between simplicity and generality. Moreover, I provide lower-level conditions for a nonparametric exchangeability condition and show that it is sufficient for Assumption 2 in Proposition 1.
@@
First, Assumption 2 can be justified by generalizing equation (2.4) of Mundlak (1978) to be $A_{i}=h\left(T^{-1}\sum_{t=1}^{T}X_{it},\nu_{i}\right)$, where $\nu_{i}\perp\left(X_{it},Z_{it}\right)$ for any $t$. Then, Assumption 2 is satisfied by taking $W_{i}=T^{-1}\sum_{t=1}^{T}X_{it}$. This is true because conditioning on $W_{i}$, any $t$-specific $X_{it}$ and $Z_{it}$ does not affect the distribution of $A_{i}$ as the residual randomness in $A_{i}$ is only driven by $\nu_{i}$. Notice that there is no need to specify the functional form of $h\left(\cdot\right)$ nor the distribution of $\nu_{i}$.
@@
Second, insights from treatment assignment models in panel data (e.g., Arkhangelsky and Imbens (2022)) can be exploited to find candidate $W_{i}$. For example, if the random vector $\left(X_{it},Z_{it}\right)^{{}^{\prime}}\in\mathbb{R}^{2}$ conditional on $A_{i}$ is a two-dimensional normal random vector i.i.d. through time with known covariance [^7] The analysis can be extended to cover the case when the covariance matrix is unknown. The required change is to expand $W_{i}$ to include the averages of the second-order polynomials of $X_{it}$ and $Z_{it}$ through time. See Chapter 7 of Hogg, McKean, and Craig (2019) for more details. and ${\bf Z}_{i}\perp A_{i}$, then $W_{i}=\left(\sum_{t=1}^{T}X_{it},\sum_{t=1}^{T}Z_{it}\right)^{{}^{\prime}}$ satisfies Assumption 2. To see this, notice that the density of $\left.\left(X_{it},Z_{it}\right)\right|A_{i}$ depend on $A_{i}$ only through the mean of $X_{it}$. Hence, by the sufficient statistic for bi-variate normal random vectors with known covariance matrix (see Chapter 7 of Hogg, McKean, and Craig (2019)), $W_{i}=\left(\sum_{t=1}^{T}X_{it},\sum_{t=1}^{T}Z_{it}\right)^{{}^{\prime}}$ suffices for Assumption 2. Furthermore, sufficient conditions for the analysis above are (i) $\left.\eta_{i}\right|A_{i}$ is a normal random vector i.i.d. through time with known variance, (ii) ${\bf Z}_{i}\perp\left(A_{i},\eta_{i}\right)$, and (iii) $g\left(\cdot\right)$ is linear in its arguments. Then, $W_{i}=\left(\sum_{t=1}^{T}X_{it},\sum_{t=1}^{T}Z_{it}\right)^{{}^{\prime}}$ satisfies Assumption 2.888Specifically, suppose $\left.\eta_{it}\right|A_{i}$ is i.i.d. normal across $t$ with mean $\mu\left(A_{i}\right)$ and known variance. Assume $X_{it}=Z_{it}^{{}^{\prime}}\gamma+A_{i}+\eta_{it}$ and ${\bf Z}_{i}\perp\left(A_{i},\eta_{i}\right)$. Let $T=2$ for simplicity. Then, define $h_{t}=x_{t}-z_{t}^{{}^{\prime}}\gamma-a$ for $t=1\text{ and }2$ and by a change of variable argument, $f_{\left.\left(X_{i1},Z_{i1},X_{i2},Z_{i2}\right)\right|A_{i}}\left(\left.x_{1},z_{1},x_{2},z_{2}\right|a\right)$ $=f_{\left.\eta_{i1},\eta_{i2}\right|A}\left(\left.h_{1},h_{2}\right|a\right)f_{Z_{i1},Z_{i2}}\left(z_{1},z_{2}\right)$ $=\left[\Pi_{t=1}^{2}f_{\left.\eta_{it}\right|A}\left(\left.h_{t}\right|a\right)\right]f_{Z_{i1},Z_{i2}}\left(z_{1},z_{2}\right).$ Using the formula of density of normal distribution for $f_{\left.\eta_{it}\right|A}\left(\left.h_{t}\right|a\right)$ and applying the Fisher-Neyman factorization theorem yield the desired result.
Third, the nonparametric exchangeability condition (Assumption 2.3) from Altonji and Matzkin (2005) can also be adapted to justify Assumption 2. A similar idea is also used by Wooldridge (2019). Notice that the original exchangeability condition of Altonji and Matzkin (2005) is incompatible with the time-varying endogeneity through the random coefficients since it concerns the density of $\eta_{it}$ given ${\bf X}_{i}$, and I present how to adapt it to the TERC model below. To grasp the idea, suppose (i) $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ [^9] To be precise, by $f_{\left.\eta_{i}\right|A_{i}}=f_{\left.\widetilde{\eta}_{i}\right|A_{i}}$ I mean $f_{\left.\eta_{i}\right|A_{i}}\left(\left.h_{i1},...,h_{iT}\right|a_{i}\right)=f_{\left.\eta_{i}\right|A_{i}}\left(\left.h_{it_{1}},...,h_{it_{T}}\right|a_{i}\right)$ for any $\left(h_{i1},...,h_{iT}\right)$ belonging to the conditional support of $\eta_{i}$ given $A_{i}=a_{i}$, where $\left(t_{1},...,t_{T}\right)$ is any permutation of $\left(1,...,T\right)$. The definition is similar for $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}}$. where $\widetilde{\eta}_{i}$ is any permutation of the vector of $\eta_{i}$ in time, (ii) $Z_{it}\perp\left(\eta_{it},A_{i}\right)$, and (iii) $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}$ is continuous in $\left({\bf X}_{i},{\bf Z}_{i}\right)$. Then, combined with Assumption 1, it is straightforward to prove that $f_{\left.A_{i}\right|{\bf X}_{i},{\bf Z}_{i}}=f_{\left.A_{i}\right|\widetilde{{\bf X}}_{i},\widetilde{{\bf Z}}_{i}}$ where $\widetilde{{\bf X}}_{i}$ and $\widetilde{{\bf Z}}_{i}$ are permutations of ${\bf X}_{i}$ and ${\bf Z}_{i}$ in time, respectively. Then, following Altonji and Matzkin (2005), I use the Weierstrass approximation theorem and the fundamental theorem of symmetric functions to obtain that the vector $W_{i}$ of elementary symmetric functions of the paired elements of $\left\{\left(X_{i1},Z_{i1}\right),...,\left(X_{iT},Z_{iT}\right)\right\}$ satisfies Assumption 2. I provide the details of the above analysis in the proof of Proposition 1. To see an example of $W_{i}$ under the nonparametric exchangeability condition, suppose $T=2$ and both $X_{it}$ and $Z_{it}$ are scalars, and define $D_{it}\coloneqq\left(X_{it},Z_{it}\right)^{{}^{\prime}}$. Then, $W_{i}=\left(\sum_{t=1}^{T}D_{it},\sum_{t=1}^{T}D_{it}D_{it}^{{}^{\prime}}\right)^{{}^{\prime}}$ satisfies Assumption 2.101010In the concrete example discussed above, the maximum order of the polynomials in the construction of $W_{i}$ depends on $T$. In general, let $\varphi_{1}\left(u\right)=\sum_{t=1}^{T}u_{t}$, $\varphi_{2}\left(u,v\right)=\sum_{t\neq s}^{T}u_{t}v_{s}$, …, $\varphi_{T}\left(u,v,...,w\right)=\sum_{t\neq s\neq...\neq k}^{T}u_{t}v_{s}...w_{k}$, where $u,v,...,w$ are generic $T\times 1$ vectors. Then, I substitute each column vector of ${\bf D}_{i}=\left({\bf X}_{i},{\bf Z}_{i}\right)$ for each of the arguments $u,v,...,w$ (repetitions included) to construct $W_{i}$. Such a $W_{i}$ satisfies Assumption 2. See Chapter II.3 of Weyl (1939) for details of the construction and its proof. Notice that parametric assumptions on the distribution of $A_{i}$ are not needed for this result.
Once the fixed effect $A_{i}$ is controlled for by $W_{i}$, I still need to control for the time-varying $\eta_{it}$ that enters the $g\left(\cdot\right)$ function and is correlated with $\beta_{it}$ via its correlation with both $A_{i}$ and $\varepsilon_{it}$. When $A_{i}$ is not present in (2.2), I may follow the method of Imbens and Newey (2009) to use the conditional CDF $V_{it}^{\text{in09}}\coloneqq F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)$ as a control for $\eta_{it}$. This is true because Theorem 1 of Imbens and Newey (2009) proves that $V_{it}^{\text{in09}}=F_{\eta_{it}}\left(\eta_{it}\right)$, the RHS of which is a one-to-one mapping of $\eta_{it}$ when $d_{X}=d_{\eta}=1$ provided (i) $Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)$, (ii) $g\left(\cdot\right)$ is strictly monotonic in $\eta_{it}$, and (iii) $F_{\eta_{it}}\left(\cdot\right)$ is a strictly increasing function. Therefore, conditional on $V_{it}^{\text{in09}}$, $X_{it}$ and $\varepsilon_{it}$ are independent. The argument above can be generalized to the setting where $d_{X}>1$ by letting
$V_{it}^{\text{in09}}$ $=\left(F_{\left.X_{it}^{\left(1\right)}\right|Z_{it}}\left(\left.X_{it}^{\left(1\right)}\right|Z_{it}\right),...,F_{\left.X_{it}^{\left(d_{X}\right)}\right|Z_{it}}\left(\left.X_{it}^{\left(d_{X}\right)}\right|Z_{it}\right)\right)^{{}^{\prime}}$
$=\left(F_{\eta_{it}^{\left(1\right)}}\left(\eta_{it}^{\left(1\right)}\right),...,F_{\eta_{it}^{\left(d_{X}\right)}}\left(\eta_{it}^{\left(d_{X}\right)}\right)\right)^{{}^{\prime}}.$
to control for the whole vector of $\eta_{it}$ coordinate by coordinate.
The challenge in applying their method to the TERC model arises from the existence of $A_{i}$ in (2.2), which is justified in many applications as agents usually know their $A_{i}$ (e.g., via repeated learning) when choosing the value of $X_{it}$. When $A_{i}$ is present in (2.2) and $d_{X}=1$, one may want to use $F_{\left.X_{it}\right|Z_{it},A_{i}}\left(\left.X_{it}\right|Z_{it},A_{i}\right)$ as the control variable for $\eta_{it}$, following a similar argument as in Imbens and Newey (2009). However, this approach is infeasible because $A_{i}$ is unknown. The idea of this paper is to replace $A_{i}$ with $W_{i}$ in the construction of $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ and show that it controls for $\eta_{it}$ under a conditional independence assumption on $Z_{it}$ and a monotonicity condition on the conditional CDF $F_{\left.\eta_{it}\right|A_{i},W_{i}}\left(\left.\eta_{it}\right|A_{i},W_{i}\right)$ in $\eta_{it}$. I summarize them in the next assumption.
@@
Suppose the following conditions are satisfied:
@@
    $F_{\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}}\left(\left.\eta_{it}^{\left(l\right)}\right|A_{i},W_{i}\right)$ is strictly increasing in $\eta_{it}^{\left(l\right)}$ for every realization of $\left(A_{i},W_{i}\right)$ on its support and each $l=1,...,d_{X}$.
Assumption 3(a) concerns the vector of exogenous IVs $Z_{it}$. Since $W_{i}$ can be viewed as summarizing all the time-invariant information about $A_{i}$ in the data, this assumption, loosely speaking, requires $\left.Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)\right|A_{i}$, which is already implied by the standard exogeneity assumption of $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$. For instance, when $d_{X}=1$ and suppose one follows Mundlak (1978) to assume $A_{i}=\left(\sum_{t=1}^{T}X_{it}\right)\alpha+\nu_{i}=W_{i}\alpha+\nu_{i}$, where $\nu_{i}$ is independent of everything else and $\alpha\neq 0$. Then, if $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it},\nu_{i}\right)$, I have $\left.Z_{it}\perp\left(\eta_{it},\varepsilon_{it}\right)\right|\left(A_{i},W_{i}\right)$ by the law of conditional independence because $W_{i}$ is a function of $A_{i}$ and $\nu_{i}$. [^11] I provide another parametric example where $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$ implies $\left.Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)\right|\left(A_{i},W_{i}\right)$. For illustration purpose, suppress subscript $i$ and suppose $T=2$, $d_{X}=d_{Z}=1$, and $W=\left(X_{1}+X_{2},Z_{1}+Z_{2}\right)^{{}^{\prime}}$ satisfies Assumption 2 (Liu, Poirier, and Shiu (2024)). Furthermore, suppose $X_{t}=Z_{t}\gamma+A+\eta_{t}$ for $t=1,2$, and $\left.\left(Z_{1},Z_{2},\varepsilon_{1},\varepsilon_{2},\eta_{1},\eta_{2}\right)\right|A\sim N\left(\mu,\Sigma\right)$ where $\Sigma=\left[\begin{array}[]{cc}\Sigma_{1}&0_{2\times 4}\\ 0_{4\times 2}&\Sigma_{2}\end{array}\right]$ and $\Sigma_{1}$ and $\Sigma_{2}$ are arbitrary positive definite $2\times 2$ and $4\times 4$ matrices, respectively. Then, formula for the conditional density function of normally distributed random vectors shows that $\left.\left(Z_{t},\varepsilon_{t},\eta_{t}\right)\right|\left(A,W\right)$ is also normally distributed with diagonal covariance matrix for $t=1,2$, satisfying Assumption 3(a). Assumption 3(a) is used to ensure that $Z_{it}$ can be excluded from the conditioning set of $F_{\left.\eta_{it}^{\left(l\right)}\right|Z_{it},A_{i},W_{i}}\left(\left.\eta_{it}^{\left(l\right)}\right|Z_{it},A_{i},W_{i}\right)$ for each $l=1,...,d_{X}$.
Assumption 3(b) is similar to condition (ii) of Theorem 1 in Imbens and Newey (2009). It is mild since it concerns a conditional CDF, which is typically strictly increasing for continuous random variables. Assumption 3(b) guarantees that fixing $\left(A_{i},V_{it},W_{i}\right)$ uniquely determines the value of $\eta_{it}$.
Assumption 3 allows me to construct a feasible control variable $V_{it}$ for $\eta_{it}$ given $A_{i}$ and $W_{i}$. To present it clearly, suppose $d_{X}=1$ and recall that $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$. Then,
@@
where the first equality holds by Assumption 2 which allows me to add $A_{i}$ to the conditioning set, the second equality holds by Assumption 1 and a change of variable argument, and the last equality holds by Assumption 3(a). Thus, $V_{it}$ uniquely determines $\eta_{it}$ given $A_{i}$ and $W_{i}$ by Assumption 3(b). Notice that $V_{it}$ is a one-to-one mapping of $\eta_{it}$ only when $A_{i}$ is also conditioned upon. In the proof of Theorem 1, I use the LIE on $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ where I first condition on $A_{i}$ to control for $\eta_{it}$ via $V_{it}$ in the inner expectation step. Then, I deal with the unknown $A_{i}$ with $W_{i}$ in the outer expectation step to prove (3.4). Finally, I use the residual variation in $X_{it}$, guaranteed by the next assumption, to identify the APE and LAR via (3.6) or (3.7).
The last assumption concerns the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$.

### v3

@@ 3.2 Assumptions
Assumption 1 (Componentwise Monotonicity).
For every $(Z_{it},A_{i})$ and each $j\in\{1,\ldots,d_{X}\}$, $g_{j}\left(Z_{it},A_{i},\eta_{it}\right)$ depends on $\eta_{it}$ only through $\eta_{it,j}$ and is strictly monotone in $\eta_{it,j}$.
The role of Assumption 1 is to deliver the invertibility/control-function step used in the identification argument. This assumption is mild when $\eta_{it}$ is scalar: it is automatically satisfied in additive triangular specifications (e.g., Newey, Powell, and Vella (1999)), and in the nonseparable scalar case it is essentially the same as the monotonicity condition in Imbens and Newey (2009, eq. (2)), except that I also allow for time-invariant heterogeneity $A_{i}$. When $d_{X}>1$, as in production-function estimation (Olley and Pakes (1996), Levinsohn and Petrin (2003), Ackerberg, Caves, and Frazer (2015)), one can further weaken the requirement to monotonicity in only one endogenous input, mirroring the identifying restriction in those classic approaches.
Assumption 1 requires justification when $\eta_{it}$ is vector-valued and each coordinate of $X_{it}$ admits one distinct coordinate of $\eta_{it}$. In the production function estimation example, this is plausible if firms make hiring and investment decisions in different units (e.g., HR and finance), each responding primarily to its own market-specific signal—labor-market conditions for hiring and financial conditions for investment—satisfying the “single-coordinate” dependence. Moreover, positive shocks in each market naturally raise the corresponding choice (more hiring when labor conditions improve, more investment when financing conditions improve), implying that $g_{j}$ is strictly monotone in $\eta_{it,j}$ for each coordinate $j$ and thus satisfying componentwise monotonicity.
Importantly, Assumption 1 places no restrictions on the dependence structure across coordinates of $\eta_{it}$, nor on the dependence between $A_{i}$ and $\eta_{it}$. Accordingly, without loss of generality, one may normalize $\eta_{it}$ so that $d_{\eta}=d_{X}$ and each coordinate of $g$ depends on the corresponding coordinate of $\eta_{it}$.
Remark 1.
Assumption 1 excludes settings where multiple components of $\eta_{it}$ enter a single coordinate of $X_{it}$, as in simultaneous supply–demand systems. Subsection 3.3 relaxes this restriction via timing assumptions, global univalence conditions, or single-index structure.
@@
Let $W_{i}\coloneqq W\left({\bf X}_{i},{\bf Z}_{i}\right)$, where $W:\mathbb{R}^{T\times\left(d_{X}+d_{Z}\right)}\mapsto\mathbb{R}^{d_{W}}$ is known. Suppose $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$ for each $t$.
Assumption 2 serves two roles. First, it makes $W_{i}$ a sufficient statistic for $A_{i}$, which is used to establish (3.9). Second, it is used to construct $V_{it}$ and establish its connection with $\eta_{it}$ by holding $A_{i}$ fixed. Under this assumption, the channel of endogeneity operating through $A_{i}$ is handled entirely via $W_{i}$, while the remaining “time-varying endogeneity through the random coefficients” is driven by the relationship between $\eta_{it}$ and $\varepsilon_{it}$ and is addressed via $V_{it}$. Assumption 2 is similar to Assumption 2.1 of Altonji and Matzkin (2005) and Assumption A2 of Liu, Poirier, and Shiu (2025). Considering the heterogeneity and endogeneity afforded in the TERC model, some restriction that isolates $A_{i}$ is typically needed.
To motivate $W_{i}$ in empirical settings, consider steady-state input policies with convex adjustment costs in the production application. In standard investment and hiring models with convex (e.g., quadratic) adjustment costs (Cooper and Haltiwanger (2006)), optimality implies convergence to a firm-specific long-run target $X_{i}^{\star}(A_{i})$ that is increasing in ability $A_{i}$. Observed inputs fluctuate around this target due to transitory shocks and gradual adjustment. Averaging $X_{it}$ over time filters out these high-frequency deviations and provides a consistent estimator of $X_{i}^{\star}(A_{i})$. Because $X_{i}^{\star}(\cdot)$ is monotone, conditioning on $\overline{X}_{i}=T^{-1}\sum_{t=1}^{T}X_{it}$ induces a control for $A_{i}$, satisfying Assumption 2.
Next, I provide three sufficient conditions for Assumption 2.
@@
First, Assumption 2 can be justified by nonparametrically generalizing equation (2.4) of Mundlak (1978) to be $A_{i}=h\left(\overline{X}_{i},\nu_{i}\right)$, where $\nu_{i}\perp\left(X_{it},Z_{it}\right)$ for all $t$. Here, the functional form of $h$ and distribution of $\nu_{i}$ can be unknown. Then, Assumption 2 is satisfied by taking $W_{i}$ to be $\overline{X}_{i}$. This is true because conditioning on $W_{i}$, any $t$-specific $X_{it}$ and $Z_{it}$ does not affect the distribution of $A_{i}$ as the residual randomness in $A_{i}$ is only driven by $\nu_{i}$.
@@
Second, insights from treatment assignment models in panel data (e.g., Arkhangelsky and Imbens (2022)) can be exploited to find candidate $W_{i}$. For example, if $(X_{it},Z_{it})$ conditional on $A_{i}$ is a two-dimensional normal random vector i.i.d. through time with known covariance and ${\bf Z}_{i}\perp A_{i}$, then the density of $(X_{it},Z_{it})$ given $A_{i}$ depends on $A_{i}$ only through $\overline{X}_{i}$. Hence, by the sufficient statistic for bi-variate normal random vectors with known covariance matrix (Hogg, McKean, and Craig (2019, ch. 7)), $W_{i}=(\overline{X}_{i},\overline{Z}_{i})$ suffices for Assumption 2.
Third, the nonparametric exchangeability condition (e.g., Altonji and Matzkin (2005)) can be adapted to justify Assumption 2. I present its details in Proposition 1 of Appendix A.2. The key idea is time exchangeability: conditional on $A_{i}$, the joint behavior of $(X_{it},Z_{it})$ is unchanged if I relabel periods, so the time order contains no information about $A_{i}$. Hence only order-invariant summaries of the panel matter for learning about $A_{i}$, and symmetric functions of the observed pairs $(X_{it},Z_{it})$ can be used as $W_{i}$ to satisfy Assumption 2.
Next, I control for the time-varying $\eta_{it}$. When $A_{i}$ is not present in (2.2) and under Assumption 1, Imbens and Newey (2009) propose
$V_{it}^{\text{IN09}}\coloneqq F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)\overset{\text{by IN09}}{=}F_{\eta_{it}}\left(\eta_{it}\right)$ (3.14)
as a control for $\eta_{it}$. The intuition is that, holding input prices $Z_{it}$ fixed, a firm choosing a higher capital $X_{it}$ must be experiencing a higher productivity shock $\eta_{it}$; otherwise the higher input choice would not be optimal. However, this argument breaks down in the TERC model because the unobserved, time-invariant component $A_{i}$ also shifts input choices: firm 1’s higher $X_{it}$ could reflect greater managerial ability rather than a larger transitory shock $\eta_{it}$. Therefore $A_{i}$ must also be conditioned on, but this makes $F_{\left.X_{it}\right|A_{i},Z_{it}}\left(\left.X_{it}\right|A_{i},Z_{it}\right)$ infeasible since $A_{i}$ is unobserved.
I proceed in two steps to generalize the argument of Imbens and Newey (2009), adapting it to the TERC model. First, to make the control for $\eta_{it}$ feasible, I replace $A_{i}$ with the observable proxy $W_{i}$ and define
$V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right).$ (3.15)
Then, I show in (A.1) that, under the next Assumption, $V_{it}$ controls for $\eta_{it}$ conditional on $A_{i}$ and $W_{i}$. Second, I further extend their analysis to allow coordinates of $g$ to depend on multiple elements of $\eta_{it}$ in Subsection 3.3.
@@
Suppose the following conditions hold:
@@
    For every $(A_{i},W_{i})$ and each $l=1,...,d_{X}$, $F_{\left.\eta_{it,l}\right|A_{i},W_{i}}(\left.\eta_{it,l}\right|A_{i},W_{i})$ is strictly increasing in $\eta_{it,l}$.
Assumption 3 enables construction of a feasible control variable $V_{it}$ for $\eta_{it}$ given $A_{i}$ and $W_{i}$. To present it clearly, suppose $d_{X}=1$ and recall that $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$. Then,
@@
where the first equality holds by Assumption 2, the second equality holds by Assumption 1 and a change of variable argument, and the last equality holds by Assumption 3(a). Thus, $V_{it}$ uniquely determines $\eta_{it}$ given $A_{i}$ and $W_{i}$ by Assumption 3(b). Note that (3.16) relies on Assumption 1; when multiple elements of $\eta_{it}$ appear in one coordinate of $X_{it}$, the one-to-one relationship does not hold in general, and additional structure or assumption is needed to recover (3.16). I discuss them in Subsection 3.3.
Assumption 3(a) is similar to condition (i) of Theorem 1 in Imbens and Newey (2009), except that I further condition on $(A_{i},W_{i})$. Since $W_{i}$ can be viewed as summarizing all the time-invariant information about $A_{i}$ in the data, Assumption 3(a), loosely speaking, requires $\left.Z_{it}\perp\left(\varepsilon_{it},\eta_{it}\right)\right|A_{i}$, which is already implied by the standard exogeneity condition $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$. Note that $Z_{it}\perp\left(A_{i},\varepsilon_{it},\eta_{it}\right)$ is stronger than Assumption 3(a) as the former rules out the possibility that $Z_{it}$ and $A_{i}$ are correlated. [^2] For instance, Bartik instruments use national industry shocks, weighted by pre-existing local industry shares. The intuition is the national shocks are plausibly “as-good-as-random” conditional on the fixed effects and controls, so the conditional exogeneity is plausible. But the baseline shares can reflect deep local traits—so unconditional exogeneity is harder to defend. Nonetheless, the unconditional exogeneity condition remains a useful benchmark since it is widely used in applied work.
Assumption 3(a) can be justified in many applications with the usual choice of IVs from the literature. For example, in labor-supply applications, $Z_{it}$ can be county minimum-wage changes or EITC expansions—policy-driven shocks that, conditional on time-invariant ability $A_{i}$ and its sufficient statistic $W_{i}$, vary independently of expected sectoral labor-demand shocks $\eta_{it}$ and labor-supply-elasticity shocks $\varepsilon_{it}$. In gasoline-demand estimation, $Z_{it}$ can be the head of household’s earned income or gas-tax changes, assumed (conditional on a household fixed effect $A_{i}$ and $W_{i}$) to be independent of anticipated near-term income changes $\eta_{it}$ (e.g., an upcoming bonus) and transitory wealth shocks $\varepsilon_{it}$.
For the empirical application of Section 5, I construct the instruments toward BLP-style (Berry, Levinsohn, and Pakes (1995)) cost shifters: leave-one-out, competitor-weighted input-price (wages and interest rates) averages at the industry–city–year level. These instruments are relevant because firms compete locally for labor and capital, so competitors’ input prices shift firm $i$’s cost environment and input choices. They are plausibly exogenous to firm $i$’s idiosyncratic productivity shocks conditional on the fixed effect and its control variable because firm $i$’s own input prices are excluded from the construction, so identification comes from competitors’ input-price variation rather than firm $i$’s unobserved shocks. Robustness checks in Appendix B provide evidence that these IVs are plausibly valid and satisfy Assumption 3(a).
Assumption 3(b) is analogous to condition (ii) of Theorem 1 in Imbens and Newey (2009). It is mild because it concerns a conditional CDF, which is typically strictly increasing for continuous random variables.
The last assumption concerns the residual variation in $X_{it}$ given $V_{it}$ and $W_{i}$, which is used to identify the APE and LAR by (3.11).